OpenStudy (anonymous):

Solve for X: 4(2x -3) + 8(x - 4) = 2(2x + 6) Please explain how I get all X's onto one side.

OpenStudy (taramgrant0543664):

The first step I would do would be to expand or in simpler terms I want rid of the brackets: 8x-12+8x-32=4x+12 Since the left side has some like terms we will collect them: 16x-44=4x+12 Next I am going to get all the values that have an x on the left side. So to get rid of the 4x on the right I will subtract it from both sides 16x-44-4x=4x+12-4x 12x-44=12 Next I will get any value that doesn't have x on the right side. So to get rid of the -44on the left I will add 44 to both sides: 12x-44+44=12+44 12x=56 To finish and get x completely by itself we have to divide each side by 12 12x/12=56/12 x=5